OpenStudy (fools101):

Need help! Please. Exponents & Logarithms How do I do this?! log4 (x - 3) + log4 (x + 3) = 2 Need help! Please. Exponents & Logarithms How do I do this?! log4 (x - 3) + log4 (x + 3) = 2 @Mathematics

OpenStudy (fools101):

Sorry didn't put this in*** Solve for x. Check for extraneous solutions...

OpenStudy (anonymous):

(x-3) *(x+3) =x^2-9 so log (base 4) x^2-9=2 is the smase the base^2 so 16=x^2-9 25=x^2 x=5 heres why x = b^y is the same as y = log(base b)x

OpenStudy (anonymous):

or x=-5 cause square root of a # could be negative

OpenStudy (fools101):

Oops sorry it not log4 it ....